For many years businesses have struggled with the rising cost of health care. But recently, the increases have slowed due to less inflation in health care prices and employees paying for a larger portion of health care benefits. A recent Mercer survey showed that 52% of U.S. employers were likely to require higher employee contributions for health care coverage. Suppose the survey was based on a sample of 400 companies. Compute the margin of error and a 95% confidence interval for the proportion of companies likely to require higher employee contributions for health care coverage. If required, round your answer to four decimal places. Round intermediate calculations to four decimal places. Margin of Error: Confidence Interval: FROM _TO_

Please explain answer using excel
Answer #1

X = np^ = 400 * 0.52 = 208

confidence interval
Input
X 208
n 400
alpha 0.05
Output
p^ 0.5200
q^ 0.4800
z-critical 1.9600
E 0.0490
lower 0.4710
upper 0.5690

formulas

confidence interval
Input
X =0.52*C5
n 400
alpha 0.05
Output
p^ =C4/C5
q^ =1-C9
z-critical =NORMSINV(1-C6/2)
E =C12*SQRT(C9*(1-C9)/C5)
lower =C9-C14
upper =C9+C14

E = margin of error = 0.0490

confidence interval from 0.471 to 0.569